# SA-ResGS: Self-Augmented Residual 3D Gaussian Splatting for Next Best View Selection

### v1

@@ 1 Introduction
Recent advances in neural rendering—particularly Neural Radiance Fields (NeRFs) (nerf) and 3D Gaussian Splatting (3DGS) (3dgs)—have significantly advanced photorealistic scene reconstruction (Yu2024MipSplatting; Niedermayr_2024_CVPR; kulhanek2024wildgaussians), enabling high-fidelity, real-time applications across diverse environments. Beyond static scene capture, these methods have spurred broader interest in tackling complex challenges, such as active view selection (Wang_2023_CVPR_EnsembleNeRF; Xiao:CVPR24:NeRFDirector; chen2024gennbv), and uncertainty quantification for next-best-view (NBV) selection (fisherrf). While pre-captured, dense-view training methods can achieve impressive reconstruction quality, in-situ (active) reconstruction—where views are selected and added progressively—remains challenging due to artifacts caused by shape-radiance ambiguity and further exacerbated by limited training views and dynamics of view addition strategy. Despite the inherent difficulty of estimating reliable uncertainty under such conditions, recent post-hoc approaches—such as Laplacian approximation-based, model-agnostic methods (fisherrf; bayesrays)—have shown promise by providing uncertainty signals without altering the rendering pipeline. However, several critical challenges remain unaddressed:
@@
While the physically grounded and uncertainty-informed selection strategy enhances overall scene coverage, it inadvertently increases sparsity in multiview overlap—since more dissimilar views are less likely to observe shared regions, thereby weakening multi-view geometric constraints. To counterbalance this contradiction without sacrificing the benefits of diverse view selection, we introduce a residual supervision mechanism that reinforces training using ground truth RGB images. The proposed SA-ResGS—particularly its residual supervision module—is fully compatible with conventional 3DGS pipelines that rely on point-cloud-based rasterization and gradient-based optimization. Beyond the original supervision, SA-ResGS additionally rasterizes color images using a selected Gaussian subset: a small fraction of the most uncertain Gaussians is combined with a majority subset (e.g., 90%) of the original visible Gaussians. This strategy mirrors the principles of residual learning and draws conceptual inspiration from Dropout (DropGaussian_CVPR2025; srivastava2014dropout) and Hard Negative Mining (xuan2020hnm; JANG2019_hnm), by amplifying supervision for Gaussians that typically receive weak gradients during backpropagation—similar to the role of skip connections in ResNet (He_2016_CVPR) in mitigating vanishing gradients. Our method thus provides additional learning signals specifically targeted at under-optimized Gaussians which are often overlooked due to their low contribution in 3DGS’s rendering mechanism, receive repeated 2D photometric supervision, resulting in more stable optimization and improved reconstruction in sparse or ambiguous regions.
@@ 2 Related Works
Next-best-view selection. NBV selection originated from the robotics community as a strategy to efficiently guide in-situ scene capture, where the goal was to incrementally select viewpoints that maximally reduce reconstruction ambiguity with minimal sensor movement (connolly1985determination; scott2003view; delmerico2018comparison). Classical NBV methods typically followed rule-based paradigms, selecting views based on geometric coverage (Dunn2009NextBV; scone; macarons), viewpoint entropy (view_entropy), or visibility heuristics (bircher2016receding; Sun_2021_CVPR). As NBV became increasingly relevant to 3D computer vision, especially under sparse-view constraints, learning-based approaches emerged to model scene-specific view policies via reinforcement learning or active learning frameworks (WANG2024_rl-based). While these data-driven methods demonstrate improved adaptability over hand-crafted rules, they often rely on task-specific reward definitions and struggle to generalize across scene types.
@@
More recent works explore representation-aware NBV strategies based on NeRF and 3DGS. Approaches such as NARUTO naruto, which learns an grid-based uncertainty field, and ActiveGAMER activegamer, which derives uncertainty from 3DGS visibility, show how representation-specific cues can guide informative view selection. Building on this direction, information-theoretic models such as FisherRF fisherrf offer a principled formulation for uncertainty-based NBV in neural fields. In our work, we build on this line (fisherrf; bayesrays; pupgs; pop_gs) by integrating physically grounded geometry priors to further stabilize early-stage view planning, particularly when minimal visual input available.
@@
Uncertainty quantification for 3DGS and neural rendering. Uncertainty quantification plays a pivotal role in active reconstruction, particularly for guiding view selection and supervision. In 3D Gaussian Splatting (3DGS), however, the high dynamic nature of primitive splitting and sensitivity to initialization often leads to unstable training, degrading the reliability of intermediate uncertainty signals. Earlier methods based on variational inference (snerf; bayesiannerf; cfnerf; lyu2024manifold) and ensemble-based estimates (ensemblenerf) enable stochastic or distributional reasoning, but they require costly model retraining or parallel inference and are typically incompatible with standard rendering pipelines. Recent post-hoc approaches (pop_gs; pupgs) such as FisherRF (fisherrf) and BayesRays (bayesrays) estimate uncertainty using Laplacian approximations without altering model structure, and have demonstrated promising results on NeRF and 3DGS variants. Complementary to geometry-based methods, image-level approaches (wang2024avs) leverage perceptual quality of current rendering results as a proxy for uncertainty. However, these methods remain strongly coupled with the density of underlying Gaussians—leading to biased uncertainty estimates in early training stages when geometry is sparse or unevenly distributed, often misinterpreting under-observed regions as confident. This overlooked bias limits the reliability of NBV guidance when it’s most needed. To mitigate this limitation, we introduce residual learning, assisted by physically grounded view selection, enabling more loosely coupled uncertainty estimation during the early-stage of view selection while emphasizing the effect of targeting high-uncertainty Gaussian focused supervision.
@@
Residual supervision in 3DGS. While accurate uncertainty estimation helps localize regions requiring stronger supervision, it alone does not guarantee that gradients effectively reach under-optimized Gaussians in the 3DGS pipeline. Residual learning, as popularized by ResNet (He_2016_CVPR), has proven effective in mitigating vanishing gradients and improving training stability through skip connections and additive refinement, yet it remains underexplored in the context of 3D Gaussian Splatting. Existing 3DGS methods mainly rely on direct photometric losses (3dgs) or external depth priors (Li_2024_CVPR_DNGaussian; Xu_ECCV_2024_MVPGS), which often fail to sufficiently supervise Gaussians with low opacity or minimal rendering contributions. Recent studies such as pixelSplat (Charatan_2024_CVPR), PAPR (Zhang_2023_NeurIPS), and PAPR-in-Motion (Peng_2024_CVPR) explicitly discuss the vanishing gradient issue and propose solutions including differentiable parameterization of Gaussians, proximity attention-based differentiable renderer, adaptive updates, and activation tuning. Despite various strategies proposed to mitigate the vanishing gradient problem, prior approaches lack an explicit mechanism for correcting weakly supervised Gaussians, which remain largely unresolved due to insufficient gradient signals. Although dropout-based approaches (DropGaussian_CVPR2025) help increase gradient diversity, they operate stochastically and do not target supervision to the most uncertain or least updated Gaussians. Our method addresses these limitations by introducing the first residual supervision strategy for 3DGS, applying uncertainty-guided rendering to intentionally amplify gradients for under-supervised Gaussians—without altering underlying rasterization process.
@@ 3 Self-Augmented Residual 3D Gaussian Splatting
The proposed SA-ResGS framework is illustrated in Fig. 1. SA-ResGS builds upon the state-of-the-art next-best-view selection method, FisherRF (fisherrf), extending it with SA-Points to support two core ideas: (1) guiding physically grounded view selection with reduced reliance on uncertainty estimation, and (2) applying residual supervision to uncertain Gaussians, mitigating the vanishing gradient problem, wherein weakly contributing Gaussians—those with minimal impact on rasterized pixel values—receive insufficient gradients during backpropagation. The physically grounded view selection is enabled by a geometrically encoded surface representation, constructed using SA-Points derived from a single training view. To construct SA-Points, we employ the 3D vision foundational model, MASt3R (mast3r), to predict dense correspondences between a given training image and a rasterized extrapolated view rendered from a nearby camera pose. The resulting 2D correspondences are then triangulated to produce 3D SA-Points.
@@ 3.1 Self-Augmented Points Generation
Given a reference image $I_{r}$ with camera pose $\mathbf{T}_{r}=[\mathbf{R}_{r}\mid\mathbf{t}_{r}]$, we render an extrapolated image $I_{e}$ from a perturbed pose $\mathbf{T}_{e}=[\mathbf{R}_{r}\mid\mathbf{t}_{r}+\Delta\mathbf{t}]$ using 3D Gaussian Splatting (3dgs). Dense correspondences $\{(\mathbf{p}_{r}^{i},\mathbf{p}_{e}^{i})\}$ between $I_{r}$ and $I_{e}$ are predicted using the pretrained MASt3R model (mast3r), which is robust to moderate viewpoint changes and capable of producing contextually meaningful matches even in the presence of minor geometric distortions. Each SA-Point $\mathbf{X}^{i}$ is triangulated from a 2D correspondence pair using the projection matrices $\mathbf{P}_{r}$ and $\mathbf{P}_{e}$ derived from COLMAP, including intrinsic matrices over each extrinsic pose $\mathbf{T}$. However, because triangulation is performed repeatedly during training—while the model is still fitting to a sparse and incomplete geometry—rasterized extrapolated images may occasionally contain rendering noise due to inaccurately placed Gaussians. To ensure reliable geometry while fully leveraging the generalization capability of MASt3R, we apply reprojection error-based filtering:
@@
This filtering step discards geometrically inconsistent points while preserving accurate SA-Points from dense, context-aware matches—even when the extrapolated image is noisier than the original training view. Compared to prior methods such as CoMapGS (jang2025comapgs) or MP-SfM (pataki2025mpsfm), our triangulation pipeline produces scale-consistent, surface-aware geometry from a single image by leveraging extrapolated viewpoints rather than requiring multiview input or monocular depth estimates. The overall steps are visualized in Fig. 2.
@@ 3.2 Physically Grounded View Selection Algorithm
where $\oplus$ denotes the element-wise XOR operation between binary vectors , and $\|\cdot\|_{1}$ is the $\ell_{1}$ norm (i.e., the number of differing entries). Here, $\mathbf{b}$ denotes a binary occupancy vector obtained by mapping voxel coordinates through a fixed random hashing function, following the spatial hashing strategy introduced in Instant-NGP (mueller2022instant). The resulting value $d_{j}\in[0,1]$ measures the proportion of voxels with inconsistent occupancy status between the currently observed volume and the candidate view. Candidate views are then ranked in descending order of their normalized Hamming distances $d_{j}$, and the top $N\%$ (e.g., $N=20$) are retained to form the physically filtered candidate set $\mathcal{C}^{\prime}$.
@@ 3.3 Uncertainty-Guided Residual Learning in 3DGS
We propose the first residual learning framework for 3DGS to address the vanishing gradient issue affecting weakly contributing Gaussians, as shown in Fig. 4. These Gaussians often receive insufficient supervision during training due to their limited impact on rasterized pixels—particularly in sparse or ambiguous regions. While ResNet (He_2016_CVPR) mitigates similar issues via skip connections, such mechanisms are infeasible in 3DGS due to the dynamic and view-dependent nature of Gaussian properties. Instead, we propose a rasterizer-agnostic strategy that enhances gradient flow by generating auxiliary renders that emphasize high-uncertainty Gaussians. These images are supervised with ground-truth RGB images, forming the basis for a residual supervision scheme detailed below.
@@
where $\lambda_{\text{full}}+\lambda_{\text{sup}}=1$, and we set both to $0.5$ in practice. We denote the losses for the full set render $I_{full}$ and the uncertainty-intensified subset render $I_{sup}$ as full loss ($\mathcal{L}_{full}$) and subset loss($\mathcal{L}_{sup}$), respectively. The uncertainty-intensified rasterization strategy is conceptually inspired by Dropout (DropGaussian_CVPR2025; srivastava2014dropout) and Hard Negative Mining (xuan2020hnm; JANG2019_hnm). Randomly sampling $\mathcal{G}_{\text{rand}}$ provides stochastic diversity, allowing weakly contributing Gaussians to be supervised when dominant ones are excluded. Meanwhile, the deterministic inclusion of $\mathcal{G}_{\text{uncertain}}$ ensures consistent gradient flow to Gaussians that are persistently under-optimized. This dual mechanism reinforces learning in uncertain or ambiguous regions without modifying the rasterization process, and complements full-image supervision to maintain global photometric fidelity.
By supervising both the full and uncertainty-intensified images, we promote stronger gradient flow toward uncertain or low-opacity Gaussians without compromising photometric quality. This strategy mirrors the role of residual skip connections in ResNet (He_2016_CVPR) (Fig. 4(b)), supporting more stable convergence and mitigating overfitting in sparse or wide-baseline training settings. It is particularly effective in the early stages of next-best-view selection, where reconstruction is sensitive to both sparsely initialized regions and supervision bias caused by overfitting to limited views.

### v2

@@ 1 Introduction
Recent advances in neural rendering—particularly Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023)—have significantly advanced photorealistic scene reconstruction (Yu et al., 2024; Niedermayr et al., 2024; Kulhanek et al., 2024), enabling high-fidelity, real-time applications across diverse environments. Beyond static scene capture, these methods have spurred broader interest in tackling complex challenges, such as active view selection (Xiao et al., 2024; Chen et al., 2024), and uncertainty quantification for next-best-view (NBV) selection (Jiang et al., 2024). While pre-captured, dense-view training methods can achieve impressive reconstruction quality, in-situ (active) reconstruction—where views are selected and added progressively—remains challenging due to artifacts caused by shape-radiance ambiguity and further exacerbated by limited training views and dynamics of view addition strategy. Despite the inherent difficulty of estimating reliable uncertainty under such conditions, recent post-hoc approaches—such as Laplacian approximation-based, model-agnostic methods (Jiang et al., 2024; Goli et al., 2024)—have shown promise by providing uncertainty signals without altering the rendering pipeline. However, several critical challenges remain unaddressed:
@@
While the physically grounded and uncertainty-informed selection strategy enhances overall scene coverage, it inadvertently increases sparsity in multiview overlap—since more dissimilar views are less likely to observe shared regions, thereby weakening multi-view geometric constraints. To counterbalance this contradiction without sacrificing the benefits of diverse view selection, we introduce a residual supervision mechanism that reinforces training using ground truth RGB images. The proposed SA-ResGS—particularly its residual supervision module—is fully compatible with conventional 3DGS pipelines that rely on point-cloud-based rasterization and gradient-based optimization. Beyond the original supervision, SA-ResGS additionally rasterizes color images using a selected Gaussian subset: a small fraction of the most uncertain Gaussians is combined with a majority subset (e.g., 90%) of the original visible Gaussians. This strategy mirrors the principles of residual learning and draws conceptual inspiration from Dropout (Park et al., 2025; Srivastava et al., 2014) and Hard Negative Mining (Xuan et al., 2020; Jang et al., 2019), by amplifying supervision for Gaussians that typically receive weak gradients during backpropagation—similar to the role of skip connections in ResNet (He et al., 2016) in mitigating vanishing gradients. Our method thus provides additional learning signals specifically targeted at under-optimized Gaussians which are often overlooked due to their low contribution in 3DGS’s rendering mechanism, receive repeated 2D photometric supervision, resulting in more stable optimization and improved reconstruction in sparse or ambiguous regions.
@@ 2 Related Works
Next-best-view selection. NBV selection originated from the robotics community as a strategy to efficiently guide in-situ scene capture, where the goal was to incrementally select viewpoints that maximally reduce reconstruction ambiguity with minimal sensor movement (Connolly, 1985; Scott et al., 2003; Delmerico et al., 2018). Classical NBV methods typically followed rule-based paradigms, selecting views based on geometric coverage (Dunn and Frahm, 2009; Guédon et al., 2022; 2023), viewpoint entropy (Vázquez et al., 2001), or visibility heuristics (Bircher et al., 2016; Sun et al., 2021). As NBV became increasingly relevant to 3D computer vision, especially under sparse-view constraints, learning-based approaches emerged to model scene-specific view policies via reinforcement learning or active learning frameworks (Wang et al., 2024a). While these data-driven methods demonstrate improved adaptability over hand-crafted rules, they often rely on task-specific reward definitions and struggle to generalize across scene types.
@@
More recent works explore representation-aware NBV strategies based on NeRF and 3DGS. Approaches such as NARUTO (Feng et al., 2024), which learns an grid-based uncertainty field, and ActiveGAMER (Chen et al., 2025), which derives uncertainty from 3DGS visibility, show how representation-specific cues can guide informative view selection. Building on this direction, information-theoretic models such as FisherRF (Jiang et al., 2024) offer a principled formulation for uncertainty-based NBV in neural fields. In our work, we build on this line (Jiang et al., 2024; Goli et al., 2024; Hanson et al., 2025; Wilson et al., 2025) by integrating physically grounded geometry priors to further stabilize early-stage view planning, particularly when minimal visual input available.
@@
Uncertainty quantification for 3DGS and neural rendering. Uncertainty quantification plays a pivotal role in active reconstruction, particularly for guiding view selection and supervision. In 3D Gaussian Splatting (3DGS), however, the high dynamic nature of primitive splitting and sensitivity to initialization often leads to unstable training, degrading the reliability of intermediate uncertainty signals. Earlier methods based on variational inference (Shen et al., 2021; Lee et al., 2025; Shen et al., 2022; Lyu et al., 2024) and ensemble-based estimates (Sünderhauf et al., 2023) enable stochastic or distributional reasoning, but they require costly model retraining or parallel inference and are typically incompatible with standard rendering pipelines. Recent post-hoc approaches (Wilson et al., 2025; Hanson et al., 2025) such as FisherRF (Jiang et al., 2024) and BayesRays (Goli et al., 2024) estimate uncertainty using Laplacian approximations without altering model structure, and have demonstrated promising results on NeRF and 3DGS variants. Complementary to geometry-based methods, image-level approaches (Wang et al., 2025b) leverage perceptual quality of current rendering results as a proxy for uncertainty. However, these methods remain strongly coupled with the density of underlying Gaussians—leading to biased uncertainty estimates in early training stages when geometry is sparse or unevenly distributed, often misinterpreting under-observed regions as confident. This overlooked bias limits the reliability of NBV guidance when it’s most needed. To mitigate this limitation, we introduce residual learning, assisted by physically grounded view selection, enabling more loosely coupled uncertainty estimation during the early-stage of view selection while emphasizing the effect of targeting high-uncertainty Gaussian focused supervision.
@@
Residual supervision in 3DGS. While accurate uncertainty estimation helps localize regions requiring stronger supervision, it alone does not guarantee that gradients effectively reach under-optimized Gaussians in the 3DGS pipeline. Residual learning, as popularized by ResNet (He et al., 2016), has proven effective in mitigating vanishing gradients and improving training stability through skip connections and additive refinement, yet it remains underexplored in the context of 3D Gaussian Splatting. Existing 3DGS methods mainly rely on direct photometric losses (Kerbl et al., 2023) or external depth priors (Li et al., 2024; Xu et al., 2024), which often fail to sufficiently supervise Gaussians with low opacity or minimal rendering contributions. Recent studies such as pixelSplat (Charatan et al., 2024), PAPR (Zhang et al., 2023), and PAPR-in-Motion (Peng et al., 2024) explicitly discuss the vanishing gradient issue and propose solutions including differentiable parameterization of Gaussians, proximity attention-based differentiable renderer, adaptive updates, and activation tuning. Despite various strategies proposed to mitigate the vanishing gradient problem, prior approaches lack an explicit mechanism for correcting weakly supervised Gaussians, which remain largely unresolved due to insufficient gradient signals. Although dropout-based approaches (Park et al., 2025) help increase gradient diversity, they operate stochastically and do not target supervision to the most uncertain or least updated Gaussians. Our method addresses these limitations by introducing the first residual supervision strategy for 3DGS, applying uncertainty-guided rendering to intentionally amplify gradients for under-supervised Gaussians—without altering underlying rasterization process.
@@ 3 Self-Augmented Residual 3D Gaussian Splatting
The proposed SA-ResGS framework is illustrated in Fig. 1. SA-ResGS builds upon the state-of-the-art next-best-view selection method, FisherRF (Jiang et al., 2024), extending it with SA-Points to support two core ideas: (1) guiding physically grounded view selection with reduced reliance on uncertainty estimation, and (2) applying residual supervision to uncertain Gaussians, mitigating the vanishing gradient problem, wherein weakly contributing Gaussians—those with minimal impact on rasterized pixel values—receive insufficient gradients during backpropagation. The physically grounded view selection is enabled by a geometrically encoded surface representation, constructed using SA-Points derived from a single training view. To construct SA-Points, we employ the 3D vision foundational model, MASt3R (Leroy et al., 2024), to predict dense correspondences between a given training image and a rasterized extrapolated view rendered from a nearby camera pose. The resulting 2D correspondences are then triangulated to produce 3D SA-Points.
@@ 3.1 Self-Augmented Points Generation
Given a reference image $I_{r}$ with camera pose $\mathbf{T}_{r}=[\mathbf{R}_{r}\mid\mathbf{t}_{r}]$, we render an extrapolated image $I_{e}$ from a perturbed pose $\mathbf{T}_{e}=[\mathbf{R}_{r}\mid\mathbf{t}_{r}+\Delta\mathbf{t}]$ using 3D Gaussian Splatting (Kerbl et al., 2023). Dense correspondences $\{(\mathbf{p}_{r}^{i},\mathbf{p}_{e}^{i})\}$ between $I_{r}$ and $I_{e}$ are predicted using the pretrained MASt3R model (Leroy et al., 2024), which is robust to moderate viewpoint changes and capable of producing contextually meaningful matches even in the presence of minor geometric distortions. Each SA-Point $\mathbf{X}^{i}$ is triangulated from a 2D correspondence pair using the projection matrices $\mathbf{P}_{r}$ and $\mathbf{P}_{e}$ derived from COLMAP, including intrinsic matrices over each extrinsic pose $\mathbf{T}$. However, because triangulation is performed repeatedly during training—while the model is still fitting to a sparse and incomplete geometry—rasterized extrapolated images may occasionally contain rendering noise due to inaccurately placed Gaussians. To ensure reliable geometry while fully leveraging the generalization capability of MASt3R, we apply reprojection error-based filtering:
@@
This filtering step discards geometrically inconsistent points while preserving accurate SA-Points from dense, context-aware matches—even when the extrapolated image is noisier than the original training view. Compared to prior methods such as CoMapGS (Jang and Pérez-Pellitero, 2025) or MP-SfM (Pataki et al., 2025), our triangulation pipeline produces scale-consistent, surface-aware geometry from a single image by leveraging extrapolated viewpoints rather than requiring multiview input or monocular depth estimates. The overall steps are visualized in Fig. 2.
@@ 3.2 Physically Grounded View Selection Algorithm
where $\oplus$ denotes the element-wise XOR operation between binary vectors , and $\|\cdot\|_{1}$ is the $\ell_{1}$ norm (i.e., the number of differing entries). Here, $\mathbf{b}$ denotes a binary occupancy vector obtained by mapping voxel coordinates through a fixed random hashing function, following the spatial hashing strategy introduced in Instant-NGP (Müller et al., 2022). The resulting value $d_{j}\in[0,1]$ measures the proportion of voxels with inconsistent occupancy status between the currently observed volume and the candidate view. Candidate views are then ranked in descending order of their normalized Hamming distances $d_{j}$, and the top $N\%$ (e.g., $N=20$) are retained to form the physically filtered candidate set $\mathcal{C}^{\prime}$.
@@ 3.3 Uncertainty-Guided Residual Learning in 3DGS
We propose the first residual learning framework for 3DGS to address the vanishing gradient issue affecting weakly contributing Gaussians, as shown in Fig. 4. These Gaussians often receive insufficient supervision during training due to their limited impact on rasterized pixels—particularly in sparse or ambiguous regions. While ResNet (He et al., 2016) mitigates similar issues via skip connections, such mechanisms are infeasible in 3DGS due to the dynamic and view-dependent nature of Gaussian properties. Instead, we propose a rasterizer-agnostic strategy that enhances gradient flow by generating auxiliary renders that emphasize high-uncertainty Gaussians. These images are supervised with ground-truth RGB images, forming the basis for a residual supervision scheme detailed below.
@@
where $\lambda_{\text{full}}+\lambda_{\text{sup}}=1$, and we set both to $0.5$ in practice. We denote the losses for the full set render $I_{full}$ and the uncertainty-intensified subset render $I_{sup}$ as full loss ($\mathcal{L}_{full}$) and subset loss($\mathcal{L}_{sup}$), respectively. The uncertainty-intensified rasterization strategy is conceptually inspired by Dropout (Park et al., 2025; Srivastava et al., 2014) and Hard Negative Mining (Xuan et al., 2020; Jang et al., 2019). Randomly sampling $\mathcal{G}_{\text{rand}}$ provides stochastic diversity, allowing weakly contributing Gaussians to be supervised when dominant ones are excluded. Meanwhile, the deterministic inclusion of $\mathcal{G}_{\text{uncertain}}$ ensures consistent gradient flow to Gaussians that are persistently under-optimized. This dual mechanism reinforces learning in uncertain or ambiguous regions without modifying the rasterization process, and complements full-image supervision to maintain global photometric fidelity.
By supervising both the full and uncertainty-intensified images, we promote stronger gradient flow toward uncertain or low-opacity Gaussians without compromising photometric quality. This strategy mirrors the role of residual skip connections in ResNet (He et al., 2016) (Fig. 4(b)), supporting more stable convergence and mitigating overfitting in sparse or wide-baseline training settings. It is particularly effective in the early stages of next-best-view selection, where reconstruction is sensitive to both sparsely initialized regions and supervision bias caused by overfitting to limited views.